The slope of the tangent line to the curve y = f(x) at (x,y) is y¹ = 10x²+5x. If (1,16) is a point on the curve

1
Expert's answer
2022-04-27T12:23:31-0400
"y'=10x^2+5x"

"y=\int(10x^2+5x)dx=\dfrac{10}{3}x^3+\dfrac{5}{2}x^2+C"

(1,16) is a point on the curve


"16=\dfrac{10}{3}(1)^3+\dfrac{5}{2}(1)^2+C"

"C=\dfrac{61}{6}"

"y=\dfrac{10}{3}x^3+\dfrac{5}{2}x^2+\dfrac{61}{6}"